A 3-phase 50Hz 2-pole 440V (L-L rms) induction motor drive has a rated speed of 2955 rpm and rated torque of 35 Nm. This motor drives a load with a torque characteristic that increases linearly with speed such that it equals the load torque at the rated speed. It draws a current of 2.1 A(rms) per phase at the rated voltage at a power factor of 0.81 (lagging). The stator resistance Rs = 2.0 ohms. At the rated operating speed, calculate the amplitude of the phase voltage (Va,braking) needed produce a regenerative braking torque that is 0.8 times the rated torque in magnitude.
Answer #1


The torque increases linearly with speed. T=KN The load torque is equal to rated torque at rated speed. 35=K (2955) K = 0.011

The stator current is, ( 440 2.1= 2+ R 0.015 R2 = 1.784 22 Substitute the values in the power equation. P=31]R, (4-) 6957.7=3